# Molecular De Novo Design through Transformer-based Reinforcement Learning

### 09 Dec 2023 v3

@@ Abstract
In this work, we introduce a method to fine-tune a Transformer-based generative model for molecular de novo design. Leveraging the superior sequence learning capacity of Transformers over Recurrent Neural Networks (RNNs), our model can generate molecular structures with desired properties effectively. In contrast to the traditional RNN-based models, our proposed method exhibits superior performance in generating compounds predicted to be active against various biological targets, capturing long-term dependencies in the molecular structure sequence. The model’s efficacy is demonstrated across numerous tasks, including generating analogues to a query structure and producing compounds with particular attributes, outperforming the baseline RNN-based methods. Our approach can be used for scaffold hopping, library expansion starting from a single molecule, and generating compounds with high predicted activity against biological targets.
@@ Introduction
Long-term Dependency Handling: Transformers utilize multi-head self-attention mechanisms which can capture long-range interactions in the data, making them particularly well-suited for modeling intricate molecular structures.
@@
Scalability: Transformers are inherently more scalable, allowing for the processing of longer sequences which is a considerable advantage in molecular design.
@@
Early de novo design algorithms were structure-based, aiming to grow ligands to fit the binding pocket of the target (Böhm, 1992; Gillet et al., 1994). However, these methods often generated molecules with poor DMPK properties and could be synthetically intractable. Ligand-based approaches were introduced to create a vast virtual library of chemical structures and then searched with a scoring function (Ruddigkeit et al., 2013; Hartenfeller et al., 2012).
@@
Molecular assembly strategies, such as string-based approaches like SMILES and SELFIES (Weininger, 1988; Krenn et al., 2020), provide an efficient representation of molecules. Graph-based methods offer an intuitive two-dimensional representation of molecular structures, with nodes and edges representing atoms and bonds, respectively (Zhou et al., 2019; Jin et al., 2018). Synthesis-based strategies, on the other hand, aim to generate only synthesizable molecules, ensuring that the design aligns with real-world applications (Bradshaw et al., 2020; Gao et al., 2022).
Various optimization algorithms have been utilized for molecular design. Genetic Algorithms (GAs) mimic natural evolutionary processes and have been applied in the context of molecule generation using both SMILES and SELFIES representations (Brown et al., 2019; Nigam et al., 2021). Bayesian optimization (BO) is another class of method that builds a surrogate for the objective function, with applications such as BOSS and ChemBO in the molecular domain (Moss et al., 2020; Korovina et al., 2020). Variational autoencoders (VAEs) offer a generative approach, mapping molecules to and from a latent space, with notable methods including SMILES-VAE and JT-VAE (Gómez-Bombarelli et al., 2018; Jin et al., 2018). Reinforcement Learning (RL) techniques, like REINVENT, have also been applied to tune models for molecule generation Olivecrona et al. (2017).
@@
In light of these developments, our approach integrates the benefits of the Transformer architecture with advanced reinforcement learning techniques, aiming to address the challenges present in current molecular de novo design methodologies.
@@
Our method first pre-trains the real 2D molecule dataset based on the transformer. Then, based on the RL paradigm, fine-tuning is performed on the molecular attributes to be optimized.
@@
The cost function , often applied to a subset of all training examples known as a batch, is minimized with respect to the network parameters . Given a predicted likelihood of the target at step , the gradient
of the prediction with respect to is used to update . This method of fitting a neural network is called back-propagation. Changing the network parameters affects not only the immediate output at time , but also influences the information flow into subsequent transformer states. This effect does not lead to problems of exploding or vanishing gradients due to the lack of recurrent connections in Transformers.
@@
Segmentation and Binary Coding of SMILES
A Simplified Molecular Input Line Entry System (SMILES) (Weininger, 2017) defines a molecule as a character sequence reflecting atoms as well as special symbols that illustrate ring opening and closure along with branching. In the majority of scenarios, SMILES are tokenized
on a single-character basis, with the exception of two-character atom types such as " " and "Br", and unique environments indicated by square brackets (e.g., , where they are processed as a single token. This approach to tokenization led to the identification of 86 tokens in the training data. Figure 3 illustrates how a chemical structure is converted to SMILES and binary-coded representations.
@@
Assume an Agent that must decide on an action to take given a particular state , where denotes the set of possible states and represents the set of potential actions for that state. The policy of an Agent associates a state to the likelihood of each action executed within. Reinforcement learning challenges are often depicted as Markov decision processes, indicating that the current state provides all essential information to inform our action choice, and no additional benefit is gained from knowing past states’ history. While this is more of an approximation than a fact for most real-life challenges, we can extend this concept to a partially observable Markov decision process where the Agent interacts with a partial environment representation. Let be the reward serving as an indicator of the effectiveness of an action taken at a certain state, and the long-term return represents the cumulative rewards collected from time to time . As molecular desirability is only meaningful for a completed SMILES, we will only consider a complete sequence’s return.
@@
For any method that necessitates a database, we exclusively use the ZINC 250K dataset 111https://www.kaggle.com/datasets/basu369victor/zinc250k. This dataset comprises approximately 250K molecules, selected from the ZINC database due to their pharmaceutical significance, manageable size, and widespread recognition. Both Screening and MolPAL conduct searches within this database. Additionally, generative models like VAEs and LSTMs are pretrained on it. Any fragments essential for JT-VAE (Jin et al., 2018), MIMOSA (Fu et al., 2021), and DST (Fu et al., 2022) are also derived from this very database.
@@
First, we compare two rule-based baselines, including:
REINVENT (Olivecrona et al., 2017).
@@
Overview: A method that employs a policy-based reinforcement learning approach to instruct RNNs to produce SMILES strings.
@@
Technical Details: It formulates the molecular design as a Markov decision process (MDP), where states represent partially generated molecules, and actions correspond to string-based manipulations. The rewards arise from properties of interest in the generated molecules.
Advantage: Adaptable and can be modified to generate other string representations, such as SELFIES.
Disadvantage: Relies heavily on the definition and design of rewards.
Graph-GA (Jensen, 2019).
@@
Overview: A genetic algorithm that manipulates molecular representations using graphs. It integrates crossover operations derived from graph matching and comprises both atom- and fragment-level mutations.
@@
Technical Details: Unlike string-based genetic algorithms, which primarily involve mutation steps, Graph-GA introduces crossover operations based on graph representations.
Advantage: Offers a richer set of operations due to its graph-based nature, potentially exploring more diverse chemical spaces.
Disadvantage: Increased complexity due to the need for graph-based operations.
SELFIES-REINVENT.
@@
Overview: An extension of REINVENT to generate SELF-referencing Embedded Strings (SELFIES).
@@
Technical Details: Like REINVENT, it utilizes a policy-based RL approach but specifically for the SELFIES representation, ensuring syntactical validity.
Advantage: Can produce molecules with fewer syntactical errors due to the nature of SELFIES.
Disadvantage: Still dependent on the definition of the reward system.
GP BO (Tripp et al., 2021).
@@
Overview: Incorporates Gaussian process Bayesian optimization, combining surrogate GP models with Graph-GA methods in inner loops.
Technical Details: While BO traditionally employs non-parametric models, GP BO leverages the GP acquisition function and integrates Graph-GA techniques for sampling.
Advantage: Balances exploration and exploitation by combining Bayesian optimization with genetic algorithms.
Disadvantage: The interplay between GP and GA might lead to higher computational costs.
STONED (Nigam et al., 2021).
@@
Overview: A modified genetic algorithm that manipulates tokens within the SELFIES strings representation.
@@
Technical Details: Unlike traditional string-based GAs, STONED directly interacts with the tokens in the SELFIES strings.
Advantage: A more direct approach that can potentially reduce invalid chemical representations.
Disadvantage: Limited to SELFIES, may not generalize to other representations.
SMILES-LSTM HC (Brown et al., 2019).
@@
Overview: An iterative learning method leveraging LSTM to comprehend the molecular distribution represented in SMILES strings.
@@
Technical Details: Uses a variant of the cross-entropy method and integrates generated high-scoring molecules into training data, subsequently fine-tuning the model.
Advantage: The iterative approach refines the generative process at each step.
Disadvantage: Convergence might be slow if the initial model is far from optimal.
SMILES-GA (Brown et al., 2019).
Overview: Genetic algorithm that defines actions based on SMILES context-free grammar.
Technical Details: Works on SMILES strings and implements genetic mutations and crossovers based on their grammar.
Advantage: Exploits the inherent structure of SMILES for effective exploration.
@@
Disadvantage: Limited to the nuances of SMILES grammar, possibly missing out on novel structures.
SynNet (Gao et al., 2022).
@@
Overview: A synthesis-based genetic algorithm that operates on binary fingerprints and decodes to synthetic pathways.
Technical Details: Focuses on the synthesis pathway, ensuring the synthesizability of generated molecules.
Advantage: Prioritizes synthesizability, ensuring generated molecules can be created in the lab.
Disadvantage: The emphasis on synthesis may limit the diversity of the molecular space explored.
DoG-Gen (Bradshaw et al., 2020).
Overview: This method is tailored to learn the distribution of synthetic pathways.
Technical Details: DoG-Gen represents synthetic pathways as Directed Acyclic Graphs (DAGs) and employs an RNN generator to model the distribution. By focusing on synthetic pathways, the method inherently emphasizes the synthesizability of molecules.
Advantages: Provides a structured approach to learning synthetic pathways.
@@
Disadvantages: Reliance on RNNs might lead to issues in capturing very long sequences if not designed effectively.
DST (Fu et al., 2022).
@@
Overview: DST stands for Differentiable Scaffolding Tree, a gradient ascent method designed for molecular optimization.
@@
Technical Details: DST abstracts molecular graphs into scaffolding trees and makes use of a graph neural network for gradient estimation. This gradient-focused approach enables fine-tuned molecular modifications based on property landscapes in the chemical space.
Advantages: Offers a more direct way to optimize molecular structures by computing gradients.
Disadvantages: The abstraction to scaffolding trees may result in loss of information.
@@
In order to evaluate both optimization capability and sample efficiency, we utilize the area under the curve (AUC) of the top-K average property value in relation to the number of oracle calls. This metric, which we refer to as AUC top-K, serves as our primary measure of performance. Distinct from the straightforward top-K average property, the AUC provides greater reward to methods that achieve high values with a reduced number of oracle calls. In this paper, we set K at 1, 10, and 100. This choice is motivated by the importance of pinpointing a limited set of unique molecular candidates for subsequent phases of development. We cap the number of oracle calls at 10,000, though we anticipate that effective methods should ideally optimize with just hundreds of calls during experimental evaluations. All AUC values reported have been min-max scaled to fit within the range [0, 1].
@@
| reinvent trans | 2.465 | 0.486 |
The evaluation results depict a thorough comparison between the REINVENT-Transformer (referred to as REINVENT-Trans) and other prominent models across multiple oracles. Overall Molecular Generation Result Performance Overview
REINVENT-Trans demonstrates its strength in molecular generation, consistently achieving top results in several oracles. For instance, the model achieved the highest performance for ‘Albuterol_Similarity’, ‘Mestranol_Similarity’, ‘QED’, ‘Scaffold_Hop’, and ‘Sitagliptin_MPO’. This suggests that the transformer’s architecture potentially excels in capturing intricate molecular patterns and relations, and effectively optimizing towards desired properties.
@@
The box plot visualizes the distribution of evaluation scores across different molecular lengths for both the reinvent_transformer method and the baseline reinvent method.
Based on the visual representation, we can derive the following observations:
1. For shorter molecular lengths, both methods exhibit similar distributions of scores.
2. As the molecular length increases, the reinvent_transformer method consistently achieves higher averge scores.
3. For longer molecular sequences, the difference in scores between the two methods becomes more pronounced. This suggests that the reinvent_transformer method is better suited for longer sequences, maintaining high evaluation scores.
4. The spread (interquartile range) of scores for the reinvent_transformer

### 10 Jan 2024 v4

@@ Abstract
In this work, we introduce a method: REINVENT-Transformer to fine-tune a Transformer-based generative model for molecular de novo design. Leveraging the superior sequence learning capacity of Transformers over Recurrent Neural Networks (RNNs), our model can generate molecular structures with desired properties effectively. In contrast to the traditional RNN-based models, our proposed method exhibits superior performance in generating compounds predicted to be active against various biological targets, capturing long-term dependencies in the molecular structure sequence. The model’s efficacy is demonstrated across numerous tasks, including generating analogues to a query structure and producing compounds with particular attributes, outperforming the baseline RNN-based methods. Our approach can be used for scaffold hopping, library expansion starting from a single molecule, and generating compounds with high predicted activity against biological targets.
@@ Introduction
Long-term Dependency Handling: Transformers utilize multi-head self-attention mechanisms, which can capture long-range interactions in the data, making them particularly well-suited for modeling intricate molecular structures.
@@
Scalability: Transformers are inherently more scalable, allowing for the processing of longer sequences, which is a considerable advantage in molecular design.
@@
Early de novo design algorithms primarily focused on structure-based methods, where the aim was to develop ligands that precisely fit the binding pocket of a target, as highlighted in works by Böhm (Böhm, 1992) and Gillet et al. (Gillet et al., 1994). These methods, while effective in certain aspects, often resulted in molecules with suboptimal drug metabolism and pharmacokinetic (DMPK) properties, and posed challenges in synthetic tractability. In contrast, ligand-based approaches, which do not rely on the 3D structure of the target, were introduced to overcome some of these limitations. They involve creating a comprehensive virtual library of chemical structures, which are then evaluated using a scoring function (Ruddigkeit et al., 2013; Hartenfeller et al., 2012). However, as noted by Blundell et al. Blundell (1996), it is important to recognize that the effectiveness of ligand-based methods compared to structure-based ones is not definitive. Both approaches have their unique advantages and limitations, and the choice between them depends on the specific requirements and context of the drug design process.
@@
Molecular assembly strategies, such as string-based approaches like SMILES and SELFIES (Weininger, 1988; Krenn et al., 2020), provide an efficient representation of molecules. Graph-based methods offer an intuitive two-dimensional representation of molecular structures, with nodes and edges representing atoms and bonds, respectively (Zhou et al., 2019; Jin et al., 2018). On the other hand, synthesis-based strategies aim to generate only synthesizable molecules, ensuring that the design aligns with real-world applications (Bradshaw et al., 2020; Gao et al., 2022a).
Various optimization algorithms have been utilized for molecular design. Genetic Algorithms (GAs) mimic natural evolutionary processes and have been applied in molecule generation using both SMILES and SELFIES representations (Brown et al., 2019; Nigam et al., 2021). Bayesian optimization (BO) is another class of method that builds a surrogate for the objective function, with applications such as BOSS and ChemBO in the molecular domain (Moss et al., 2020; Korovina et al., 2020). Variational autoencoders (VAEs) offer a generative approach, mapping molecules to and from a latent space, with notable methods including SMILES-VAE and JT-VAE (Gómez-Bombarelli et al., 2018; Jin et al., 2018). Reinforcement Learning (RL) techniques, like REINVENT, have also been applied to tune models for molecule generation Olivecrona et al. (2017).
@@
The evolution of molecular design methodologies has progressively addressed various challenges and limitations. The transition from RNN-based methods to more advanced generative models underscores a quest for improved handling of complex chemical structure representations and optimization. While RNNs brought significant progress, their inherent difficulty in capturing long-term dependencies in sequential data has been a notable shortcoming. This gap is precisely where the Transformer architecture, introduced by Vaswani et al. (Vaswani et al., 2023), brings its strengths to the fore. Its self-attention mechanism allows for a more nuanced and effective handling of sequence data, a critical aspect in molecular design where long-range interactions within molecules play a pivotal role.
Parallelly, the field has seen the integration of reinforcement learning (RL) for fine-tuning generative models, as evidenced in the work by Jaques et al. (Jaques et al., 2017). RL’s ability to iteratively improve models based on a feedback loop aligns well with the demands of molecular design, where continuous refinement based on molecular properties is essential. The combination of RL with generative models has been shown to enhance the ability to navigate the vast chemical space more effectively, achieving better results in molecule generation.
Therefore, in light of these advancements and limitations, we propose an approach that integrates the Transformer architecture with advanced RL techniques. This proposal is underpinned by the Transformer’s superior handling of sequential data and the iterative refinement capability of RL. By merging these two powerful technologies, we aim to address the existing challenges in molecular de novo design, such as the need for better sequence representation and optimization. This integration promises to enhance the effectiveness and efficiency of molecular generation processes, moving closer to achieving more sophisticated and automated molecular design systems.
@@
Our method, named as REINVENT-Transformer, first pre-trains the real 2D molecule dataset based on the transformer. Then, based on the RL paradigm, fine-tuning is performed on the molecular attributes to be optimized.
Preliminaries
In this study, our focus is confined to versatile single-objective molecular optimization techniques that are pertinent to the design of small organic molecules. These molecules possess scalar properties that are significant in the context of therapeutic development. The molecular design challenge at hand can be formally described as an optimization task:
Here, represents a molecular structure, while is the expansive domain known as chemical space, encompassing all potential molecular candidates. The extent of is overwhelmingly vast, for instance, around Bohacek et al. (1996). It’s presupposed that we have access to the actual value of a targeted property, symbolized by , in which an oracle, , functions as an opaque mechanism. This oracle assesses specific chemical or biological attributes of a molecule and yields the real property as a scalar value. It’s important to note that the oracles do not provide an analytical form or derivatives of the properties. The most feasible oracles—either experimental procedures or high-fidelity simulations—often come with considerable expense. Therefore, an algorithm that can efficiently optimize the oracle within a feasible resource allocation is crucial. Such an algorithm would be a key component in the automation of molecular design, contributing significantly to advanced automated chemical design (ACD) Goldman et al. (2022) or function-driven autonomous synthesis Gao et al. (2022b).
@@
Training data Overview: Segmentation and Binary Coding of SMILES
A Simplified Molecular Input Line Entry System (SMILES) (Weininger, 2017) defines a molecule as a character sequence reflecting atoms as well as special symbols that illustrate ring opening and closure along with branching. In the majority of scenarios, SMILES are tokenized on a single-character basis, with the exception of two-character atom types such as " " and "Br", and unique environments indicated by square brackets (e.g., , where they are processed as a single token. This approach to tokenization led to the identification of 86 tokens in the training data.
@@
A transformer block is a parameterized function class . If then where
The notation indicates we take the softmax (defined in Equation 7) over the -dimensional vector indexed by . The LayerNorm function Ba et al. (2016) is defined for by
The set of parameters, denoted by , comprises the elements of the weight matrices and the LayerNorm parameters and , as specified on the right-hand side. The input represents a set of entities, each characterized by attributes (typically, though not exclusively, sequences of -dimensional vectors of length ). It is important to note that the output retains the same format as the input . A transformer is an amalgamation of distinct transformer blocks, each equipped with unique parameters: . Key hyperparameters in a transformer include , and , with typical configurations being . While the initial research suggested , more recent studies tend to employ a greater number of these blocks.
@@
The cost function , often applied to a subset of all training examples known as a batch, is minimized with respect to the network parameters . Given a predicted likelihood of the target at step , the gradient of the cost function with respect to is used to update . This method of fitting a neural network is called back-propagation. Changing the network parameters affects not only the immediate output at time , but also influences the information flow into subsequent transformer states.
@@
Agent Decision-Making and Markov Decision Processes Assume an Agent that must decide on an action to take given a particular state , where denotes the set of possible states and represents the set of potential actions for that state. The policy of an Agent associates a state to the likelihood of each action executed within. Reinforcement learning challenges are often depicted as Markov decision processes, indicating that the current state provides all essential information to inform our action choice, and no additional benefit is gained from knowing past states’ history. While this is more of an approximation than a fact for most real-life challenges, we can extend this concept to a partially observable Markov decision process where the Agent interacts with a partial environment representation. Let be the reward serving as an indicator of the effectiveness of an action taken at a certain state, and the long-term return represents the cumulative rewards collected from time to time (Sutton & Barto, 1999). As molecular desirability is only meaningful for a completed SMILES, we will only consider a complete sequence’s return.
@@
Negative Log-Likelihood (NLL) and Loss Function To assess the likelihood of sequence generation by the agent, we use the Negative Log-Likelihood (NLL). The NLL is calculated as follows:
This measure is crucial in understanding the generative model’s performance (Blaschke et al., 2020). The augmented likelihood and loss function are then computed to adjust the agent’s generation process:
Scoring Functions for Molecular Sequences REINVENT-Transformer utilizes scoring functions to evaluate and guide the generation of molecular sequences. These functions are formulated as either a weighted product or a weighted sum:
This scoring approach is designed to balance various molecular properties during the generation process, facilitating the production of molecules with desired characteristics (Blaschke et al., 2020).
@@
For any method that necessitates a database, we exclusively use the ZINC 250K dataset Irwin & Shoichet (2005). This dataset comprises approximately 250K molecules, selected from the ZINC database due to their pharmaceutical significance, manageable size, and widespread recognition. Both Screening Ahmed et al. (2018) and MolPAL Graff et al. (2021) conduct searches within this database. Additionally, generative models like VAEs Kingma & Welling (2014) and LSTMs Yu et al. (2019) are pretrained on it. Any fragments essential for JT-VAE (Jin et al., 2018), MIMOSA (Fu et al., 2021), and DST (Fu et al., 2022) are also derived from this very database.
@@
First, we compare two rule-based baselines, shown in the Table 1
| Method | Overview | Technical Details | Advantage | Disadvantage |
REINVENT |
@@
A method employing a policy-based reinforcement learning approach to instruct RNNs to produce SMILES strings. |
@@
Formulates molecular design as a Markov decision process with states representing partially generated molecules and actions as string manipulations. Rewards based on properties of interest. |
Adaptable to generate other string representations like SELFIES. |
Heavily reliant on the design of rewards. |
Graph-GA |
@@
A genetic algorithm that manipulates molecular representations using graphs, with graph matching and atom/fragment mutations. |
@@
Introduces crossover operations based on graph representations, unlike string-based genetic algorithms. |
Offers a richer set of operations for exploring diverse chemical spaces. |
Increased complexity due to graph-based operations. |
SELFIES-REINVENT |
@@
An extension of REINVENT for generating SELF-referencing Embedded Strings (SELFIES). |
@@
Uses a policy-based RL approach for SELFIES representation, ensuring syntactical validity. |
Produces molecules with fewer syntactical errors. |
Still dependent on reward system definition. |
GP BO |
@@
Combines Gaussian process Bayesian optimization with Graph-GA methods. |
Leverages GP acquisition function integrated with Graph-GA techniques for sampling. |
Balances exploration and exploitation effectively. |
Higher computational costs due to GP and GA interplay. |
STONED |
@@
A modified genetic algorithm that manipulates tokens within SELFIES strings. |
@@
Interacts directly with tokens in SELFIES strings, differing from traditional string-based GAs. |
Direct approach potentially reduces invalid chemical representations. |
Limited to SELFIES, may not generalize to other representations. |
SMILES-LSTM HC |
@@
Iterative learning method using LSTM to understand molecular distribution in SMILES strings. |
@@
Employs a variant of the cross-entropy method, fine-tuning the model with high-scoring molecules. |
Iteratively refines the generative process. |
Slow convergence if initial model is suboptimal. |
SMILES-GA |
Genetic algorithm based on SMILES context-free grammar. |
Implements genetic mutations and crossovers based on SMILES grammar. |
Exploits SMILES structure for effective exploration. |
@@
Confined to SMILES grammar nuances, potentially missing novel structures. |
SynNet |
@@
Synthesis-based genetic algorithm operating on binary fingerprints and decoding to synthetic pathways. |
Focuses on synthesizability of generated molecules. |
Prioritizes synthesizability, ensuring lab producibility of molecules. |
Limited diversity in molecular space exploration due to synthesis emphasis. |
DoG-Gen |
Tailored to learn the distribution of synthetic pathways. |
Represents synthetic pathways as DAGs, using an RNN generator for modeling. Emphasizes synthesizability. |
Structured approach to learning synthetic pathways. |
@@
Issues in capturing very long sequences with RNNs if not designed effectively. |
DST |
@@
Differentiable Scaffolding Tree method for molecular optimization using gradient ascent. |
@@
Abstracts molecular graphs into scaffolding trees, using a graph neural network for gradient estimation. |
Direct optimization of molecular structures through gradient computation. |
Possible loss of information due to abstraction to scaffolding trees. |
@@
In order to evaluate both optimization capability and sample efficiency, we utilize the area under the curve (AUC) of the top- average property value in relation to the number of oracle calls. This metric, which we refer to as AUC top-K, is formally defined as follows:
Given a sequence of molecules generated by a method, and an oracle function that returns the property value of a molecule, the top- average property value at any point in the sequence is given by:
where is the -th highest property value molecule among the first molecules.
The AUC top-K is then the area under the curve when plotting the top- average property value against the number of oracle calls up to molecule , for to . This is calculated as:
We set at 1, 10, and 100, capping the number of oracle calls at 10,000. All AUC values reported are min-max scaled to the range [0, 1].
Recall (Sensitivity): Traditionally, recall is the proportion of actual positives correctly identified. In our context, it is the proportion of molecules with desirable properties (as judged by the oracle) that the method successfully identifies from the total ’N’ molecules deemed desirable by the oracle.
Precision (Positive Predictive Value): Precision is the proportion of predicted positives that are true positives. Here, it is the proportion of molecules identified by the method as having desirable properties that are indeed validated by the oracle, out of the ’M’ molecules selected by the method.
@@
| REINVENT-Trans | 2.465 | 0.486 |
| Model | SMILES | Score | Number |
| REINVENT-Transformer | Cc1csc(NC(=O)c2ccc(N3CCCC3=O)cc2)n1 | 0.9479 | 1656 |
| REINVENT-Transformer | COc1cc(NC(=O)c2cnn(C)c2)cc(Cl)c1Cl | 0.9477 | 1875 |
| REINVENT-Transformer | Cc1ncsc1CNC(=O)c1cc(C(F)(F)F)cn1C | 0.9475 | 1873 |
| REINVENT-Transformer | Cc1cc(C(F)(F)F)nn1CC(=O)Nc1ccc(C#N)cc1 | 0.9474 | 466 |
| REINVENT | CS(=O)(=S)c1ccc(C(=O)Nc2ccc(F)cc2)cc1 | 0.9481 | 6853 |
| REINVENT | Cc1ccc(C(=O)Nc2c(F)cc(F)cc2C(=O)N(C)C)o1 | 0.9481 | 5825 |
| REINVENT | Cc1ccc(C(=O)Nc2ccc(S(C)(=O)=O)c(F)c2)cc1 | 0.9481 | 4525 |
| REINVENT | Cc1ccc(S(C)(=O)=O)cc1C(=O)Nc1ccc(F)cc1 | 0.9481 | 4605 |
The evaluation results depict a thorough comparison between the REINVENT-Transformer (referred to as REINVENT-Trans) and other prominent models across multiple oracles. Randomly selected SMILES generated by different models can be seen in Table 4. And the corresponding chemical structures are shown in figure 2
Overall Molecular Generation Result Performance Overview
REINVENT-Transformer demonstrates its strength in molecular generation, consistently achieving top results in several properties. For instance, the model achieved the highest performance for ‘Albuterol_Similarity’, ‘Mestranol_Similarity’, ‘QED’, ‘Scaffold_Hop’, and ‘Sitagliptin_MPO’. This suggests that the transformer’s architecture potentially excels in capturing intricate molecular patterns and relations, and effectively optimizing towards desired properties.
@@
The box plot visualizes the distribution of evaluation scores across different molecular lengths for both the REINVENT-Transformer method and the baseline REINVENT method.
Based on the figure 3, we can derive the following observations:
1. In general, REINVENT-Transformer will generate longer average length of molecules than REINVENT.
2. The REINVENT-Transformer method consistently achieves higher average scores.
3. The spread (interquartile range) of scores for the REINVENT-Transformer